OpenStudy (anonymous):

f(x)=x/(x+1) and g(x)=x^2-1 How do I find this function (f+g)(x)

OpenStudy (campbell_st):

well \[(f + g)(x) = f(x) + g(x)\] so you have \[(f+g)(x) = \frac{x}{x + 1} + x^2 - 1\]

OpenStudy (anonymous):

Yeah I got that far but then how do I solve that problem? @campbell_st

OpenStudy (campbell_st):

ok... so use a common denominator \[(f + g)(x) = \frac{x}{x +1} + \frac{(x^2 -1)(x +1)}{x + 1}\] which becomes \[(f + g)(x) = \frac{x}{x +1} + \frac{x^3 + x^2 - x - 1}{x + 1}\] so I'll leave you to collect like terms in the numerator to simplify the equation

OpenStudy (anonymous):

(f+g)(x) = x^3+x^2-2x-1/(x+1)

OpenStudy (campbell_st):

well if you have x - x what do you have... you need to just check the numerator